# Rotatable Antenna Enabled Wireless Communication: Modeling and Optimization

### v2

@@ IV-B Two-Stage Algorithm
s.t. $\rho_{k}\bar{P}_{k}\|\hat{\mathbf{f}}^{T}\hat{\mathbf{a}}_{k}\|^{2}\geq\omega,\;\forall k,$ (49b)
$\hat{\mathbf{f}}_{3(n-1)+1:3n}^{T}\mathbf{e}_{1}\geq\operatorname{cos}\theta_{\mathrm{max}},\;\forall n,$ (49c)
$\|\hat{\mathbf{f}}_{3(n-1)+1:3n}\|^{2}=1,\;\forall n,$ (49d)
where $\rho_{k}$ can be initially set as $\rho_{k}=1-\rho_{\mathrm{ZF},k}$ based on (48) with $\bm{\Theta}=\mathbf{0}$, serving as a weight factor for user $k$ to be balanced for maximizing the channel power gains of all users, and constraints (49c) and (49d) are equivalent to (38c) and (38d), respectively. Note that $\|\hat{\mathbf{f}}^{T}\hat{\mathbf{a}}_{k}\|^{2}=\hat{\mathbf{f}}^{T}\mathbf{A}_{k}\hat{\mathbf{f}}=\mathrm{Tr}(\mathbf{A}_{k}\hat{\mathbf{f}}\hat{\mathbf{f}}^{T})$ with $\mathbf{A}_{k}=\hat{\mathbf{a}}_{k}\hat{\mathbf{a}}_{k}^{H}$. Define $\hat{\mathbf{F}}=\hat{\mathbf{f}}\hat{\mathbf{f}}^{T}$, which needs to satisfy $\hat{\mathbf{F}}\succeq\mathbf{0}$ and $\mathrm{rank(\hat{\mathbf{F}})}=1$. Since the rank-one constraint is non-convex, we apply semidefinite relaxation (SDR) technique to relax this constraint. As a result, problem (P9) is transformed into
@@
s.t. $\rho_{k}\bar{P}_{k}\mathrm{Tr}(\mathbf{A}_{k}\hat{\mathbf{F}})\geq\omega,\;\forall k,$ (50b)
$\hat{\mathbf{F}}_{3(n-1)+1,3(n-1)+1}\geq\operatorname{cos}^{2}\theta_{\mathrm{max}},\;\forall n,$ (50c)
$\mathrm{Tr}(\hat{\mathbf{F}}_{3(n-1)+1:3n,3(n-1)+1:3n})=1,\forall n,$ (50d)
$\hat{\mathbf{F}}\succeq\mathbf{0}.$ (50e)
@@
As problem (P10) is a convex semidefinite program (SDP), it can be optimally solved by the CVX solver with a complexity order of $\mathcal{O}((3N)^{3.5})$ [39], and we represent the optimal solution to $\hat{\mathbf{F}}$ by $\hat{\mathbf{F}}^{\star}$.
@@
Since problem (P10) may not lead to a rank-one solution of $\hat{\mathbf{F}}$, the optimal objective value of problem (P10) serves as an upper bound of problem (P9). Thus, a rank-one approximation on $\hat{\mathbf{F}}^{\star}$ should be executed as an additional step to construct a feasible solution to problem (P9). If $\hat{\mathbf{F}}^{\star}$ is rank-one, we have $\hat{\mathbf{F}}^{\star}=\hat{\mathbf{f}}^{\star}\hat{\mathbf{f}}^{\star T}$, and $\hat{\mathbf{f}}^{\star}$ will be a feasible and optimal solution to problem (P9). On the other hand, if the rank of $\hat{\mathbf{F}}^{\star}$ is larger than one, we define $\tilde{\mathbf{f}}=\sqrt{\lambda_{\mathrm{max}}}\mathbf{m}_{\mathrm{max}}$ with $\lambda_{\mathrm{max}}$ and $\mathbf{m}_{\mathrm{max}}$ denoting the maximum eigenvalue and its corresponding eigenvector obtained through eigenvalue decomposition of $\hat{\mathbf{F}}^{\star}$, respectively, as our candidate solution to problem (P9) since the best rank-one approximation to $\hat{\mathbf{F}}^{\star}$ is given by $\tilde{\mathbf{F}}^{\star}=\lambda_{\mathrm{max}}\mathbf{m}_{\mathrm{max}}\mathbf{m}_{\mathrm{max}}^{T}$ [39].
@@
2) Second Stage: After obtaining the stacked pointing vector $\hat{\mathbf{f}}^{\star}$ in the first stage and letting $\mathbf{f}_{n}=\hat{\mathbf{f}}_{3(n-1)+1:3n}^{\star},\;\forall n$, we reconstruct the RA deflection angles similar to (45) and the channel conditions according to (8)–(10). Then, the corresponding ZF beamforming can be readily calculated by (47).
@@
Remark 2: Although the two-stage algorithm may experience some performance loss by limiting $p=1$, it has lower complexity without the need for iteration. In contrast, the AO algorithm, while having higher computational complexity, is applicable to any value of $p$. It is worth noting that both the AO and two-stage algorithms can also be applied to the single-user setup with arbitrary $Q$ scatterers by replacing the MMSE/ZF receive beamforming of (34)/(47) with MRC beamforming. Nevertheless, for the single-user setup with $Q=0$, since the AO and two-stage algorithms can only achieve suboptimal solutions and result in higher computational complexity, they are much less efficient than the optimal closed-form solution derived in (18) of Section III-A.
@@ V-A Single-User System
First, we consider a single-user system under free-space propagation, where the distance between the center of the array and the user is set to $r=$ 15 m. In this subsection, we consider only the cosine gain pattern (i.e., $p=\frac{1}{2}$) to validate the performance analysis in Section III-B. The received signal power $P_{R}$ at the BS, which is proportional to the resultant SNR due to the relationship $P_{R}=\sigma^{2}\gamma$, is considered as the performance metric.
@@
Figure 6: Maximum received signal power versus the number of antennas $N_{x}$ for the ULA-based single-user system.
@@
To compare the directional gains of RA and fixed-antenna systems, Fig. 6 plots the maximum received signal power versus the number of antennas $N_{x}$ for a ULA-based system. The received signal power is obtained by averaging over various user directions. The asymptotic values given in (24) and (25) are also shown in the figure. It is first observed that the closed-form SNR derived in (27) matches perfectly with the exact value calculated by (20), which validates the correctness of Theorem 1 and Lemma 1. Additionally, for a small to moderate number of antennas, the received signal powers of both the RA and fixed-antenna systems increase linearly with $N_{x}$, which is in accordance with Remark 1. However, as $N_{x}$ further increases, it is observed that due to the limited deflection angle range for antenna boresight adjustment, the received signal powers of both RA and fixed-antenna systems increase slowly and eventually approach to their asymptotic values. Furthermore, due to the additional spatial DoFs induced by the antenna boresight adjustment for improving the array directional gain, RA system reaches its asymptotic limit later and achieves up to 1.43 dB gain over the fixed-antenna system, which corroborates the accuracy of analytical result in (26) as $10\mathrm{log}_{10}\frac{\lim_{N_{x}\to\infty}\tilde{\gamma}}{\lim_{N_{x}\to\infty}\tilde{\gamma}_{\mathrm{fixed}}}=10\mathrm{log}_{10}\left(\frac{\pi}{6}+\operatorname{cos}\frac{\pi}{6}\right)\approx$ 1.43 dB. Based on this result, we can also infer that with a larger antenna directivity factor $p$, the RA system will exhibit a more pronounced array gain over the fixed-antenna system due to the higher directional gain in the boresight direction of each antenna towards the user.
Figure 7: Maximum received signal power versus the number of antennas $N$ for the square UPA-based single-user system.
In Fig. 7, we plot the received signal power versus the number of antennas for a square UPA-based system, where $N_{x}=N_{y}$ and the position of the user is given by $[0,0,r]^{T}$. First, it can be observed that for the RA system, both the analytical lower and upper bounds in (31) are quite tight and accurate to the exact value of the received signal power calculated via (20). With a small to moderate number of antennas, the received signal powers of both the RA and fixed-antenna systems increase linearly with $N$, which is in accordance with our approximation in Lemma 2. As the number of antennas further increases, the performance advantage of the RA system over the fixed-antenna system becomes more significant. This is because the beam direction will deviate from the boresight direction of the antenna located at the edge of the array in the fixed-antenna system, which degrades the array gain. Instead, this deviation can be effectively circumvented by adjusting the antenna boresight direction to align with the user direction as closely as possible.
@@
Figure 8: Maximum received signal power versus the azimuth angle of the user $\phi$ for the UPA-based single-user system with $N_{x}$ = $N_{y}$ = 201.
@@
Fig. 8 shows the received signal power versus the azimuth angle of the user for a square UPA, where the zenith angle of the user is given by $\psi=\frac{\pi}{2}$. As $\phi$ increases from 0 to $\frac{\pi}{3}$ or decreases from 0 to $-\frac{\pi}{3}$, the received signal power of the fixed-antenna system drastically decreases. This is due to the fact that the array directional gain pattern of the fixed-antenna system is fixed and the radiation power only focuses on the region directly in front of the array as shown in Fig. 3. In contrast, the gain pattern of the RA array can be flexibly reconfigured to improve the directional gain in the user direction, even when the user directly faces one of the two ends of the array. It can be observed that the RA array helps the BS achieve more stable and uniform received signal power over the entire angular region as compared to the fixed-antenna system. The above results indicate that RA array has the potential to uniformly enhance communication coverage performance in its front half-space.
@@ V-B Multi-User System
Next, we consider a multi-user system under the multi-path channel model with $K=$ 4 users and $Q=$ 3 scatterers, and their geometric relationships are shown in Fig. 9. Specifically, four users lie evenly on a half circle centered at the BS array with a radius of $d_{1}=$ 50 m, and three scatterers are randomly located within three disk regions with a radius of $d_{3}=$ 2 m, which lie evenly on a half circle centered at the center of the BS array with a radius of $d_{2}=$ 25 m. A square UPA-based system with $N_{x}=N_{y}=$ 9 is considered. In the following, we present simulation results by averaging over 100 independent channel realizations, and the max-min achievable rate (which monotonically increases with the SINR) among all the $K$ users is considered as the performance metric, which is given by
$C=\min_{k}\;\mathrm{log}_{2}\left(1+\gamma_{k}\right)=\mathrm{log}_{2}\left(1+\min_{k}\;\gamma_{k}\right).$ (51)
To validate the performance advantages of our proposed RA-enabled system, we consider the following three benchmark schemes for comparison:
@@
  • Random orientation design: In this scheme, the deflection angles of each RA, i.e., $\{\theta_{\mathrm{z},n}\}$ and $\{\theta_{\mathrm{a},n}\}$, are randomly generated following a uniform distribution within $[0,\theta_{\mathrm{max}}]$ and $[0,2\pi]$, respectively, and the MMSE receive beamforming is applied at the BS.
@@
  • Array-wise orientation adjustment: In this scheme, we adjust the orientation of the entire antenna array instead of that of each antenna element. To achieve the max-min achievable rate, by applying the MMSE receive beamforming, the optimal zenith and azimuth angles of the antenna array are obtained by using exhaustive search within $[0,\theta_{\mathrm{max}}]$ and $[0,2\pi]$, respectively.
@@
  • Without orientation adjustment (i.e., fixed-antenna system): In this scheme, the orientations of all RAs are fixed as their reference orientations, i.e., $\bm{\Theta}=\mathbf{0}$, and the MMSE receive beamforming is applied at the BS.
Figure 9: Simulation setup of the multi-user and multi-path channel case.
@@
In Fig. 10, we compare the max-min achievable rates versus the user transmit power $P$ for different optimization algorithms. First, it is observed that the AO algorithm with MMSE receive beamforming obtains the highest max-min achievable rate and achieves up to 3 dB gain over the two-stage algorithm. As discussed in Remark 2, this gain gap is expected since the lack of iteration and the assumption of $p=1$ in the two-stage algorithm will inevitably cause performance loss. Second, by optimizing the deflection angles of all RAs to maximize the minimum weighted channel power gain with much lower computational complexity than the AO algorithm, the two-stage algorithm still achieves up to 2.5 dB gain over the fixed-antenna system without orientation/boresight optimization. The above results indicate the different tradeoffs between performance and complexity offered by the two proposed algorithms. Furthermore, since the ZF receiver enhances the noise power, the AO algorithm with MMSE beamforming always outperforms that with ZF beamforming.
@@
Figure 10: Max-min achievable rates obtained by different optimization algorithms versus the user transmit power $P$.
@@
Fig. 11 shows the max-min achievable rates of different schemes versus the maximum zenith angle $\theta_{\mathrm{max}}$. Several interesting observations are made as follows. First, as the maximum zenith angle $\theta_{\mathrm{max}}$ increases, the proposed RA system gains more DoFs and flexibilities to balance the array directional gain over the multi-path channels, thus leading to a further increase in its max-min achievable rate. Second, the proposed RA system always outperforms both the array-wise orientation adjustment counterpart and the fixed-antenna system. This is because neither of the latter two systems can independently adjust the orientation of each antenna to reconfigure the directional gain pattern of the entire array, and the radiation power is distributed only directly in front of the array. Third, since the RAs with random orientations can statistically radiate power in any direction of the BS to serve spatially-distributed users, it can achieve a higher max-min achievable rate than the fixed-antenna system. However, when $\theta_{\mathrm{max}}\geq\frac{3\pi}{10}$, the max-min achievable rate of the random orientation design scheme slightly declines with $\theta_{\mathrm{max}}$. This result highlights the importance of antenna orientation/boresight optimization in an RA system, since the random orientations will lead to an unordered array directional gain pattern and inevitable performance loss when $\theta_{\mathrm{max}}$ is large enough. Last but not least, it is interesting to note that the growth speed of the max-min achievable rate obtained by our proposed RA-enabled system substantially increases when $\theta_{\mathrm{max}}\leq\frac{\pi}{10}$, which indicates that even with a small deflection angle range for RA orientation/boresight adjustment, the proposed RA-enable system with optimized deflection angles can achieve significant performance improvement.
@@
Figure 11: Max-min achievable rates of different schemes versus the maximum zenith angle $\theta_{\mathrm{max}}$.
@@
Figure 12: Max-min achievable rates of different systems versus the antenna directivity factor $p$.
@@
Fig. 12 shows the max-min achievable rates of different schemes versus the antenna directivity factor $p$. It is observed that the max-min achievable rate of the proposed RA-enabled system increases with the directivity factor $p$. This can be explained by the fact that with a larger $p$, the directional gain in the boresight direction of the antenna becomes larger and the main lobe becomes narrower, which is more advantageous for our RA-enabled system to enhance directional gains in multiple user directions by adjusting the orientations/boresights of RAs, thus resulting in a larger max-min achievable rate. In contrast, the max-min achievable rates of the array-wise orientation adjustment scheme and the fixed-antenna system decrease with $p$ when $p\geq 1.5$. The reason for this is that with a larger directivity factor $p$, the radiation power of both schemes will be more concentrated in the region directly in front of the array. As a result, the directional gains for users deviating from the main direction of the array will become weaker, thus resulting in a lower max-min achievable rate. Additionally, although the random orientation design scheme can disperse the radiation power of the array in multiple specific directions, it is significantly inferior to the proposed RA-enabled system since it fails to strategically allocate the antenna resources to fairly improve the communication performance of all users. The above results highlight the necessity of our proposed RA-enabled system for increasing channel capacity, especially in situations where the antennas have strong directivity, i.e., the main lobes are narrow.

### v4

@@ IV-B Two-Stage Algorithm
s.t. $\rho_{k}\bar{P}_{k}\|\mathbf{f}^{T}\mathbf{m}_{k}\|^{2}\geq\omega,\;\forall k,$ (53b)
$\mathbf{f}_{3(n-1)+1:3n}^{T}\mathbf{e}_{3}\geq\operatorname{cos}\theta_{\mathrm{max}},\;\forall n,$ (53c)
$\|\mathbf{f}_{3(n-1)+1:3n}\|^{2}=1,\;\forall n,$ (53d)
where $\rho_{k}$ can be initially set as $\rho_{k}=1-\rho_{\mathrm{ZF},k}$ based on (52) with $\vec{\mathbf{f}}_{n}=\mathbf{e}_{3},\;\forall n$, serving as a weight factor for user $k$’s channel power gain, and constraints (53c) and (53d) are equivalent to (42c) and (42d), respectively. Note that $\|\mathbf{f}^{T}\mathbf{m}_{k}\|^{2}=\mathbf{f}^{T}\mathbf{M}_{k}\mathbf{f}=\mathrm{Tr}(\mathbf{M}_{k}\mathbf{f}\mathbf{f}^{T})$ with $\mathbf{M}_{k}=\mathbf{m}_{k}\mathbf{m}_{k}^{H}$. Define $\bar{\mathbf{F}}=\mathbf{f}\mathbf{f}^{T}$, which needs to satisfy $\bar{\mathbf{F}}\succeq\mathbf{0}$ and $\mathrm{rank(\bar{\mathbf{F}})}=1$. Since the rank-one constraint is non-convex, we apply semidefinite relaxation (SDR) technique to relax this constraint. As a result, problem (P9) is transformed into
@@
s.t. $\rho_{k}\bar{P}_{k}\mathrm{Tr}(\mathbf{M}_{k}\bar{\mathbf{F}})\geq\omega,\;\forall k,$ (54b)
$\bar{\mathbf{F}}_{3(n-1)+1,3(n-1)+1}\geq\operatorname{cos}^{2}\theta_{\mathrm{max}},\;\forall n,$ (54c)
$\mathrm{Tr}(\bar{\mathbf{F}}_{3(n-1)+1:3n,3(n-1)+1:3n})=1,\forall n,$ (54d)
$\bar{\mathbf{F}}\succeq\mathbf{0}.$ (54e)
@@
As problem (P10) is a convex semidefinite program (SDP), it can be optimally solved by the CVX solver with a complexity order of $\mathcal{O}((3N)^{3.5})$ [46], and we represent the optimal solution of problem (P10) as $\left\{\omega^{\star},\bar{\mathbf{F}}^{\star}\right\}$.
@@
Since problem (P10) may not lead to a rank-one solution of $\bar{\mathbf{F}}$, the optimal objective value of problem (P10) serves as an upper bound of problem (P9). Thus, a rank-one approximation on $\bar{\mathbf{F}}^{\star}$ should be executed as an additional step to construct a feasible solution to problem (P9). If $\bar{\mathbf{F}}^{\star}$ is rank-one, we have $\bar{\mathbf{F}}^{\star}=\mathbf{f}^{\star}\mathbf{f}^{\star T}$, and $\mathbf{f}^{\star}$ will be a feasible and optimal solution to problem (P9). On the other hand, if the rank of $\bar{\mathbf{F}}^{\star}$ is larger than one, we define $\bar{\mathbf{f}}=\sqrt{\lambda_{\mathrm{max}}}\bm{\upsilon}_{\mathrm{max}}$ with $\lambda_{\mathrm{max}}$ and $\bm{\upsilon}_{\mathrm{max}}$ denoting the maximum eigenvalue and its corresponding eigenvector obtained through eigenvalue decomposition of $\bar{\mathbf{F}}^{\star}$, respectively, as our candidate solution to problem (P9) since the best rank-one approximation to $\bar{\mathbf{F}}^{\star}$ is given by $\mathbf{f}^{\star}=\lambda_{\mathrm{max}}\bm{\upsilon}_{\mathrm{max}}\bm{\upsilon}_{\mathrm{max}}^{T}$ [46].
@@
2) Second Stage: After obtaining the stacked pointing vector $\mathbf{f}^{\star}$ in the first stage and letting $\vec{\mathbf{f}}_{n}=\mathbf{f}_{3(n-1)+1:3n}^{\star},\;\forall n$, we reconstruct the channels according to (8)–(10). Then, the corresponding ZF beamforming can be calculated by (50).
The computational complexity of the proposed two-stage algorithm depends on the SDP and singular value decomposition in the first stage, and the calculation of the ZF beamforming in the second stage. Thus, the complexity order of the two-stage algorithm is given by $\mathcal{O}\left(KN^{3}+(3N)^{3}+(3N)^{3.5}\right)$. Since the two-stage algorithm only needs to solve problem (P10) and calculate the beamforming in (50) for one time, it requires lower computational complexity than the AO algorithm, especially when the number of antennas is large or the convergence accuracy of the AO algorithm is strict.
@@
Remark 3: Although the two-stage algorithm may experience some performance loss by setting $p=1$, it has lower complexity without the need for iteration. In contrast, the AO algorithm, while having higher computational complexity, is applicable to any value of $p$. It is worth noting that both the AO and two-stage algorithms can also be applied to the single-user setup with arbitrary $Q$ scatterer clusters by replacing the MMSE/ZF receive beamforming of (37)/(50) with MRC beamforming. Nevertheless, for the single-user setup with $Q=0$, since the AO and two-stage algorithms can only achieve suboptimal solutions and result in higher computational complexity, they are much less efficient than the optimal closed-form solution derived in (20) of Section III-A. On the other hand, although the original problem can be transformed into a pointing vector optimization problem by substituting the MMSE receive beamforming of (37) into (13), it is found that the objective function becomes even more complicated due to the presence of matrix inversion, which is generally difficult to handle and thus not considered in this paper.
@@ V-A Single-User System
First, we consider a single-user system under free-space propagation, where the distance between the center of the array and the user is set to $r=$ 15 m. In this subsection, we consider the cosine gain pattern (i.e., $p=\frac{1}{2}$) to validate the performance analysis in Section III-B. The received signal power $P_{R}$ at the BS, which is proportional to the resultant SNR due to the relationship $P_{R}=\sigma^{2}\gamma$, is considered as the performance metric.
@@
Figure 5: Maximum received signal power versus the number of antennas $N_{x}$ for the ULA system with $\phi=\frac{5\pi}{12}$ and $\psi=\frac{\pi}{2}$.
@@
To compare the directional gains of RA and fixed-antenna systems, Fig. 5 plots the maximum received signal power versus the number of antennas $N_{x}$ for a ULA system. The asymptotic values given in (27) and (28) are also shown in the figure. It is first observed that the closed-form SNR derived in (30) matches perfectly with the exact value calculated by (23), which validates the correctness of Theorem 1 and Lemma 1. Additionally, for a small to moderate number of antennas, the received signal powers of both the RA and fixed-antenna systems increase linearly with $N_{x}$, which is in accordance with Remark 2. However, as $N_{x}$ further increases, it is observed that the received signal powers of both RA and fixed-antenna systems eventually approach their asymptotic values. Meanwhile, the RA system reaches its asymptotic limit later and achieves up to 1.43 dB gain over the fixed-antenna system, which corroborates the accuracy of analytical result in (29) as $10\mathrm{log}_{10}\frac{\lim_{N_{x}\to\infty}\tilde{\gamma}}{\lim_{N_{x}\to\infty}\tilde{\gamma}_{\mathrm{fixed}}}=10\mathrm{log}_{10}\left(\frac{\pi}{6}+\operatorname{cos}\frac{\pi}{6}\right)\approx$ 1.43 dB. Furthermore, when $N_{x}\leq 100$, the RA system achieves up to 5 dB gain over the conventional fixed-antenna system. This indicates that the proposed RA architecture can achieve a high array gain due to the additional spatial DoFs induced by the antenna boresight rotation.
@@
Figure 6: Maximum received signal power versus the antenna directivity factor $p$ for the UPA system with $N_{x}=N_{y}=4$ and $\psi=\frac{\pi}{2}$.
In Fig. 6, we plot the received signal power versus the antenna directivity factor $p$ for the UPA system. It is observed that the received signal power of the RA system increases as $p$ increases. This is expected since a higher directivity (i.e., larger $p$) results in greater radiation power in the boresight direction, leading to a higher array gain focused on the user direction when the boresight of each antenna is aligned as closely as possible with the user direction. Conversely, the received signal power of the fixed-antenna system first increases and then decreases with $p$. This is because, with a relatively low antenna directivity (e.g., $p\leq 4$), the main lobe is sufficiently broad to encompass the user direction, allowing the fixed antenna to offer enhanced directional gain compared to the isotropic antenna. However, as $p$ continues to increase, the main lobe becomes narrower, eventually excluding the user direction from its effective coverage. Consequently, the directional gain in the user direction decreases with the further increase of $p$. Moreover, it is observed that the received signal power of the RA system is consistently higher than that of the fixed-antenna system regardless of $p$, highlighting the performance advantage of the RA system, which benefits from the radiation power focusing capability enabled by orientation/boresight adjustments.
Figure 7: Maximum received signal power versus the azimuth angle of the user $\phi$ for the UPA system with $N_{x}=N_{y}=4$ and $\psi=\frac{\pi}{2}$.
@@
Fig. 7 shows the received signal power versus the azimuth angle of the user $\phi$ for the UPA system, where the zenith angle of the user is given by $\psi=\frac{\pi}{2}$. As $\phi$ increases from 0 to $\frac{\pi}{2}$ or decreases from 0 to $-\frac{\pi}{2}$, the received signal power of the fixed-antenna system drastically decreases. This is due to the fact that the array directional gain pattern of the fixed-antenna system is fixed and the radiation power only focuses on the region directly in front of the array. In contrast, the RA array enables the BS to maintain more stable and uniform received signal power over the entire angular region, particularly when $\phi$ lies within the interval $[-\frac{\pi}{6},\frac{\pi}{6}]$, where each RA can adjust its boresight within the allowable rotational range to precisely align with the user direction. In addition, the RA system achieves much higher received signal power than the fixed-antenna system even when the user direction deviates significantly from the array’s main direction, i.e., $\phi\to-\frac{\pi}{2}$ or $\frac{\pi}{2}$. These advantages stem from the RA array’s ability to flexibly reconfigure its directional gain pattern to enhance the directional gain in the user direction. The above results indicate that RA array has the potential to uniformly enhance communication coverage performance in its front half-space.
@@ V-B Multi-User System
Next, we consider a multi-user system under the multipath channel model with $K=4$ users and $Q=8$ scatterer clusters. Specifically, four users are uniformly distributed in four distinct directions in front of the BS. The distance of each user to the BS is independently and randomly selected from a uniform distribution within $[30,50]$ m. Around these users, eight scatterer clusters are randomly distributed. In the following, we present simulation results by averaging over 500 independent channel realizations. Meanwhile, the max-min effective communication rate (which monotonically increases with the SINR) among all the $K$ users is considered as the performance metric, which is given by
$C$ $=\min_{k}\;(1-\varrho)\mathrm{log}_{2}\left(1+\gamma_{k}\right)$
$=(1-\varrho)\mathrm{log}_{2}\left(1+\min_{k}\;\gamma_{k}\right),$ (55)
where the weighting coefficient $\varrho$ with $10\leq\varrho<$ 1 is introduced to account for the system overhead. In our simulations, the weighting coefficients for the RA and fixed-antenna systems are set as $\varrho_{\mathrm{RA}}=0.05$ and $\varrho_{\mathrm{fixed}}=0$, respectively.
In Fig. 8, we illustrate the convergence behavior of the proposed AO algorithm (i.e., Algorithm 1) with different numbers of RAs. It is observed that the max-min effective communication rate increases over iterations and converges within six iterations. This indicates that the proposed AO algorithm converges quickly and demonstrates its effectiveness. To further validate the performance advantages of our proposed RA system, we consider the following four benchmark schemes for comparison:
@@
  • Random orientation design: In this scheme, the orientation of each RA is randomly generated within the rotational ranges given by (4), and the MMSE receive beamforming is applied at the BS.
@@
  • Array-wise orientation adjustment: In this scheme, we adjust the orientation of the entire antenna array instead of that of each antenna element. By applying the MMSE receive beamforming, the optimal orientation of the antenna array is obtained by using exhaustive search.
@@
  • Fixed orientation design: In this scheme, the orientations of all RAs are fixed at their reference orientations, i.e., $\vec{\mathbf{f}}_{n}=\mathbf{e}_{3},\;\forall n$, and the MMSE receive beamforming is applied at the BS.
  • Baseline with isotropic antennas: In this scheme, the antenna elements in the array are isotropic, i.e., $p=0$ and the radiation energy is evenly distributed in the front half-space of the antennas, and the MMSE receive beamforming is applied at the BS.
Figure 8: Convergence behavior of the proposed AO algorithm.
@@
In Fig. 9, we compare the max-min effective communication rates versus the user transmit power $P$ for different optimization algorithms. First, it is observed that the AO algorithm with MMSE receive beamforming obtains the highest max-min effective communication rate and achieves up to 2.5 dB gain over the two-stage algorithm. As discussed in Remark 3, this gain gap is expected since the lack of iteration and the assumption of $p=1$ in the two-stage algorithm inevitably cause performance loss. Second, by optimizing the pointing vectors of all RAs to maximize the minimum weighted channel power gain with much lower computational complexity than the AO algorithm, the two-stage algorithm still achieves up to 5 dB gain over the fixed-antenna system without orientation/boresight optimization. The above results indicate the different trade-offs between performance and complexity offered by the two proposed algorithms. Specifically, the AO algorithm can achieve better performance due to its iterative optimization process, while the two-stage algorithm offers competitive performance at significantly lower computational overhead. Furthermore, since the ZF receiver enhances the noise power, the AO algorithm with MMSE beamforming always outperforms that with ZF beamforming.
@@
Figure 9: Max-min effective communication rates obtained by different optimization algorithms versus the user transmit power $P$.
@@
Fig. 10 shows the max-min effective communication rates of different schemes versus the maximum zenith angle $\theta_{\mathrm{max}}$. Several interesting observations are made as follows. First, as $\theta_{\mathrm{max}}$ increases, the proposed RA system gains more DoFs and flexibilities to balance the array directional gain over the multipath channels, thus leading to a further increase in its max-min effective communication rate. Second, the proposed RA system always outperforms both the array-wise orientation adjustment counterpart and the fixed-antenna system. This is because neither of the latter two systems can independently adjust the orientation/boresight of each individual antenna to reconfigure the directional gain pattern of the entire array. Third, since the RAs with random orientations can statistically radiate power in any direction of the BS to serve spatially-distributed users, it can achieve a higher max-min effective communication rate than the fixed-antenna system. However, when $\theta_{\mathrm{max}}\geq\frac{3\pi}{10}$, the max-min effective communication rate of the random orientation design scheme declines with $\theta_{\mathrm{max}}$. This result highlights the importance of antenna orientation/boresight optimization in an RA system, since the random orientations will lead to an unordered array directional gain pattern and inevitable performance loss when $\theta_{\mathrm{max}}$ becomes large. Last but not least, it is interesting to note that the growth rate of the max-min effective communication rate obtained by our proposed RA system substantially increases when $\theta_{\mathrm{max}}\leq\frac{\pi}{10}$, which indicates that even with a small rotational range for RA orientation/boresight adjustment, the proposed RA system with optimized pointing directions can achieve significant performance improvement.
@@
Figure 10: Max-min effective communication rates of different schemes versus the maximum zenith angle $\theta_{\mathrm{max}}$.
@@
Figure 11: Max-min effective communication rates of different systems versus the antenna directivity factor $p$.
@@
Fig. 11 shows the max-min effective communication rates of different schemes versus the antenna directivity factor $p$. It is observed that the max-min effective communication rate of the proposed RA system increases with the directivity factor $p$. This is attributed to the fact that a larger $p$ leads to a higher directional gain in the antenna’s boresight direction and a narrower main lobe, which is more advantageous for the proposed RA system. Consequently, the RA system with larger $p$ can more effectively enhance directional gains across multiple user directions, thereby achieving a larger max-min effective communication rate. In contrast, the max-min effective communication rates of the array-wise orientation adjustment scheme and the fixed-antenna system decrease with $p$ when $p\geq 1$. The reason for this is that with a larger directivity factor $p$, the radiation power of both schemes will be more concentrated in the region directly in front of the array. As a result, the directional gains for users deviating from the main direction of the array will become weaker, thus resulting in a lower max-min effective communication rate. Additionally, although the random orientation design scheme can disperse the radiation power of the array in multiple specific directions, it is significantly inferior to the proposed RA system since it fails to strategically allocate the antenna resources to fairly improve the communication performance of all users. The above results highlight the necessity of our proposed RA system for increasing channel capacity, especially in situations where the antennas have strong directivity, i.e., their main lobes are narrow.